You are here: Foswiki>Commissioning Web>Com20120204 (05 Feb 2012, JuanGuerra)Edit Attach

UT 20120204 AGw#3 at LFBG Verification

Observer: DMiller, JCarlos, JHill (LBTO)
Telescope Operator: DHuerta (LBTO)
Telescope Support: JUrban(LBTO)
Instrument Support: DOfficer (LBTO)

Summary

Details

1:50 Opened chamber

2:00 David center on BS9107 by changing IE, CA, but they were pretty c

2:10 preset BS9107 ACTIVE Failed, no star in acquisition image. This is because the current on-axis position value are off by 7 mm so ~14". The active preset and the acquire preset had PA that were 120 degrees apart. Thus, the star rotated out of the field of view of the guide camera when the rotator moved

Guide Probe Center of Rotation

2:30 preset BS9107 ACTIVE, PA=180. This had guide probe position at x=0, y = 611.2

2:40 John move the guide probe by hand (setxy) to x = -7.2 y = 609.2, Bob Meeks' value of on-axis

4:50 Dave rotation agw from 0 to 360. Doug takes 80 second image on guider /OldRepository/AGW_Data/20120204/guider_center/left_guider000002.fits

  • Find center on image at pixel (128, 148) (binned x2)
  • Moved probe to x=-4.2, y =609.2

3:00 Dave rotation agw from 360 to 0. Doug takes 80 second image on guider /OldRepository/AGW_Data/20120204/guider_center/left_guider000004.fits

  • Find center on image at pixel (174, 149) (binned x2)
  • Moved probe to x=-1.2, y =610.2

3:07 Dave rotation agw from 0 to 360. Doug takes 80 second image on guider /OldRepository/AGW_Data/20120204/guider_center/left_guider000004.fits

  • Find center on image at pixel (224, 165) (binned x2)
  • OOPS, guide image is binned by 2. Must divide hotspot number by 2 - old hotspot is 274,298 unbinned, or 137,149 binned

  • Moved probe back to x=-6.9, y =609.2 (or -469 1002245 in encoder counts)

3:10 Wind is up to peak value of 18 m/s frown, sad smile

3:15 Dave rotation agw from 0 to 360. Doug takes 80 second image on guider /OldRepository/AGW_Data/20120204/guider_center/left_guider000004.fits

  • Find center on image at pixel (132, 149)
  • John inserts the corresponding values into oacontrol.cfg file.

Using the example from 20111101, we should add the -469 to the homeoffset value, and we should add 3300 microns to the ooffset value.
homeoffset=129531 and ooffset=418800

At the same time John subtracted the homeoffset adjustment from the lim_f and lim_n values. Since Thomas had change homeoffset to 130000 in November, John assumed that the theta limits needed that same big adjustment.

Guider Hotspot at 70 elevation

3:40 collimated at 70 elevation

  • getHotspot left 274,298
  • setHotspot left 270, 298 OK
  • setHotspot left 265, 298 OK
  • setHotspot left 260, 298 OK image 62
  • setHotspot left 255, 298 OK image 64
  • setHotspot left 250, 298 OK image 66
  • setHotspot left 245, 298 OK image 68
  • setHotspot left 240, 298 OK image 70
  • setHotspot left 235, 298 OK image 72
  • setHotspot left 230, 298 OK image 74 Vignetting on right upper
  • setHotspot left 270, 298 OK image 77
  • setHotspot left 290, 298 OK image 79 Vignetting left
  • setHotspot left 270, 298 OK image 81
  • setHotspot left 270, 280 OK image 84 OK
  • setHotspot left 270, 270 OK image 87
  • setHotspot left 270, 265 OK image 89 Vignetting on bottom
  • setHotspot left 270, 298 OK image 91
  • setHotspot left 270, 310 OK image 93 A little Vignetting on the top
  • setHotspot left 270, 315 OK image 95 Too much vignetting on the top
  • setHotspot left 270, 298 OK image 77

4:00 Done collecting data

  • Hotspot should be at x = 260, y = 288

Guide Probe Center of Rotation Again

The restart procedure is: Stop GCS, stopAGW -u 3, stop oacserver, check that oacserver is gone, start oacserver, startAGW -u 3, start GCSL, wait for GCS to home probes.

4:05 John has new center position and limits in oaccontrol. David stop AGw3, restart oaccontrol, GCS. Probe would not move to center position, and wouldn't move away from parked position at all. Apparently John's assumptions about the limits were wrong. See correct adjustments below.

4:30 John removed limit changes and only include changes to center of rotation position.

4:40 David spin rotator from 0 to 360. Doug take 90 second image /OldRepository/AGW_Data/20120204/guider_center/left_guider000007.fits

  • Find center of rotation way of to the left of the image by 200+ pixels. John decides his x correction had the wrong sign. Later figured out that and x-flip had been inserted after the November adjustment so the example had been wrong. Now the correction is +469 to make homeoffset=130469.

4:44 David stop AGw3, restart oaccontrol, GCS.

4:47 David acquire star. Then preset with same position angle. Star is still off to the side of the acquisition image - apparently farther. We are confused.

4:54 David spin rotator from 0 to 360. Doug take 90 second image /OldRepository/AGW_Data/20120204/guider_center/left_guider000008.fits

  • Find center of rotation way of to the left of the image by 200+ pixels. Did change not take effect

5:00 David make sure oaccontrol is dead, restart, and try again.

5:20 David spin rotator from 0 to 360. Doug take 90 second image /OldRepository/AGW_Data/20120204/guider_center/left_guider000009.fits

  • Find center of rotation way of to the left of the image by 400+ pixels. Wrong direction

5:35 John changed the delta position last made by a factor of 3 and opposite sign. homeoffset=127200 John stop AGw3, start stop aoccontrol, GCS.

  • Find an even larger circle!! How did we make the sign wrong 3 times in a row? We decide that when we changed stars (ie position on the sky) the apparent position of the center rotated and our analysis was incorrect.

6:00 John change value to -100, in between the first two values, image 12 homeoffset=129900

  • Find center of rotation just on the right edge of the guider image. Better. Change to +100 homeoffset=130100

6:40 Restart all, acquire, recenter, active preset, image 13

  • Find center of rotation 1/4 the width of the guide image from the right side. Better.
  • Center measured at 429, 274 (unbinned)
  • Change to +300 homeoffset=130300

6:50 Restart all, acquire, recenter, active preset, image 14

  • Find center of rotation
  • Center at same place
  • Change was not read properly ????

7:10 Trouble with preset. Could not move probe. Tried to re-home but could not. Reboot agw-control computer

7:33 acquire, center star, active preset

7:40 Restart all, acquire, recenter, active preset, image 15

  • Find center of rotation at (335, 275)
  • Change to +429 homeoffset=130429

7:53 Restart all, acquire, recenter, active preset, image 16

  • Find center of rotation at (274, 272) Great!! - exactly what we wanted in X.

Guider Hotspot at 34 elevation

8:00

  • getHotspot left 274,298
  • setHotspot left 220 298 image 146: half Vignetted on the right
  • setHotspot left 310 298 image 156: half Vignetted on the left
  • setHotspot left 265 250 image 164: half Vignetted on the bottom
  • setHotspot left 265 320 image 167: half Vignetted on the top

  • Hotspot should be at x = 265, y = 290

  • Put this in LUCI_L.cfg

Guide Probe Center of Rotation Again

8:20

  • Change to +442 homeoffset=130442
  • Change ooffset to 418220 from 418800

8:35 Restart all, acquire, recenter, active preset, image 18

  • Find center of rotation at (266, 293)
  • Now set to new (just determined above) hotspot values below

9:00 Update /home/telescope/TCS/Configuration/GCS/LUCI_L.cfg

  • guidecam_hotspot_x 265.0
  • guidecam_hotspot_y 290.0
  • WFScam_hotspot_x 386.0
  • WFScam_hotspot_y 243.0
  • WFS_lenslet_edge_x 346.0
  • WFS_lenslet_edge_y 213.0

9:05 David restart GCS. Failed, could not initialize guide camera. stopAGW, startAGW, home and try again. Turned out that Doug made a typo in LUCI_L.cfg

Later: After consulting with Jesper and Thomas, John updated the theta limits lim_f and lim_n by subtracting 442 from the values Thomas had set in November.

This concludes the probe alignment to the center of rotation.

Pisces

09:40 Switch over to PISCES at RFBG for some tests there. Images taken with the ND filter holder in place. Defocusing a center bright star and we can not see a round defocused image.

The star position in Pisces after a preset is out of the detector, it is necesary to apply an offset of:x=5, y=1 to put the star in the center. Calculate the rotaiton center with and offset in x=7 y=1. Images are:
ROT=0 pisces.20120204_0046.fits
ROT=60 pisces.20120204_0047.fits
ROT=120 pisces.20120204_0049.fits
ROT=180 pisces.20120204_0050.fits
ROT=240 pisces.20120204_0051.fits
ROT=300 pisces.20120204_0052.fits
ROT=360 pisces.20120204_0053.fits

Aquisition of star images in the four corners of the detector:

upper right pisces.201204_0056.fits
upper left pisces.201204_0057.fits
lower left pisces.201204_0058.fits
lower right pisces.201204_0059.fits
center pisces.201204_0060.fits

To determine the alignment defocuse of a bright star is used. The star is centered in the detector and defocused, this shows a pattern that doen't looks like a defocused images.

A set of defocused images of 10 seconds exposure were taken in the four corners. The filter is clear. The list is:

center pisces.201204_0073.fits
lower left pisces.201204_0074.fits
upper right pisces.201204_0075.fits
upper left pisces.201204_0076.fits
lower right pisces.201204_0079.fits
center pisces.201204_0080.fits

A serie of defocused images in the uppoer left corner were taken, the image is moving inside.

upper left edge pisces.201204_0082.fits zero
upper left edge pisces.201204_0083.fits one arcsec in
upper left edge pisces.201204_0085.fits two arcsec in

Debugging the new observing modes for AO observations.

12:49:38 Preset MODE=ACE. success RUNAO failed due an error; " 12:49:38. Camera Lens too cold T=5.7, must be above 6.0". The close loop command is not send.

13:06:56. MODE=ADAPTIVEACE_TRACK . preset success. MODE=ADAPTIVEACE_TRACK, ROT=PARALLACTIC, preset success and the ROTAT is in hold position.

-- DougMiller - 04 Feb 2012
Topic revision: r10 - 05 Feb 2012, JuanGuerra
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback